How many milliliters of orange puree must be mixed with 300 ml of 30% orange to obtain 45% orange concentrate?
Answer by stanbon(75887)

How many milliliters of orange puree must be mixed with 300 ml of 30% orange to obtain 45% orange concentrate?
-----------------
x + 0.30*300 = 0.45(x+300)
----
Multiply thru by 100 to get:
100x + 30*300 = 45x + 45*300
55x = 15*300
x = (3/11)300
---
x = 81.82 ml (amt of 100% orange puree needed)
================
Cheers,
Stan H.
================